### v1

# Shock-capturing particle hydrodynamics with reproducing kernels

Abstract
We present and explore a new shock-capturing particle hydrodynamics approach. We enhance a commonly used SPH discretization with Roe’s approximate Riemann solver where we use slope-limited linear reconstruction in the dissipative terms of the solver. All gradients of the method are calculated with linearly reproducing kernels that are constructed to enforce the two lowest order consistency relations so that constant and linear functions can be recovered to machine precision. We scrutinize the method with a set of challenging 3D benchmark problems ranging from shocks over instabilities to Schulz-Rinne-type vorticity creating shocks. In all of them we find excellent agreement with analytic/reference solutions.
keywords:
hydrodynamics – methods: numerical – instabilities – shock waves – software: simulations
pagerange: Shock-capturing particle hydrodynamics with reproducing kernelsLABEL:lastpagepubyear: 2024
1 Introduction
The Smoothed Particle Hydrodynamics (SPH) method has originally been developed (Lucy, 1977; Monaghan, 1977) to solve astrophysical gas dynamics problems. In an SPH simulation, the particle distribution automatically adapts to the dynamics of the gas flow and this has major advantages when, for example, modeling the gravitational collapse of gas clouds that condense into denser filaments and finally form stars. The natural geometric adaptivity has also advantages in many other contexts where challenging geometries are involved, e.g in simulating dam breaks (e.g. Xu & Yu, 2021), or fracture processes (e.g. Rahimi & Moutsanidis, 2022), see Monaghan (2012) for a broad range of further applications. Since particles automatically follow the gas flow, this also implies that vacuum is simply modeled by the absence of computational particles, while Eulerian methods need to model vacuum as a low-density background gas. This can cost substantial computational resources even though one is not interested (and noting is physically going on) in empty space. As an (admittedly extreme, but astrophysically relevant) example, we show in Fig. 1 the tidal disruption of two stars by a massive black hole that is located at the coordinate origin (the simulation is discussed in more detail in Rosswog (2023)). As the stars pass by the black hole, they are ripped apart by the hole’s tidal forces. Clearly, one is only interested in the fate of the gas from the disrupted stars which initially cover only a minute fraction of $10^{-9}$ of the space that is shown at $t=371$ h. So in a corresponding Eulerian simulation one would need to waste essentially all of the computational resources to simulate the uninteresting, empty space. Another major advantage of SPH is that can be formulated in a way so that Nature’s conservation laws are enforced by construction (Monaghan, 1992, 2005; Rosswog, 2009; Springel, 2010; Price, 2012; Rosswog, 2015b). Having nature’s conservation laws built into a simulation gives one some confidence that the simulated system behaves similarly to what nature does.
Simulating gas flows with particles is still a relatively young field compared to Eulerian gas dynamics and much development is still ongoing. SPH has initially been criticized for a number of issues, but on essentially all of them major improvements have been achieved. For example, SPH has often been criticized for being overly dissipative. As derived from a Lagrangian, however, SPH contains zero dissipation and the criticism goes back to the early days of SPH when simple artificial dissipation schemes were used in which dissipation was always switched on, whether needed or not. Much effort has been spent in recent years to cure this issue, e.g. by via time-dependent dissipation parameters that only reach substantial values when needed, but not otherwise (Morris & Monaghan, 1997; Rosswog et al., 2000; Cullen & Dehnen, 2010; Rosswog, 2015a, 2020a) or by using slope-limited reconstruction in the artificial dissipation terms (Frontiere et al., 2017; Rosswog, 2020b; Rosswog & Diener, 2021; Rosswog et al., 2023; Sandnes et al., 2024). The latter approach has been shown to be very effective even if large constant dissipation parameters are used. For example, weakly triggered Kelvin-Helmholtz instabilities, that do not grow when standard, constant dissipation parameters are used, show a healthy growth close to the theoretical expectation when reconstruction is applied, see Fig. 20 in Rosswog (2020b).
As an alternative to artificial dissipation, one can also implement Riemann solvers into SPH (Inutsuka, 2002; Parshikov & Medin, 2002; Cha et al., 2010; Murante et al., 2011; Sirotkin & Yoh, 2013; Puri & Ramachandran, 2014; Meng et al., 2021) to produce sufficient entropy in shocks. But also with this approach, one needs to make sure that not too much unwanted dissipation is introduced. For example, simply treating particle pairs as Riemann problems where the left and the right state are given by the particle properties leads to very dissipative hydrodynamics unless other measures such as limiters or a reconstruction techniques similar to Finite Volume schemes are applied. Often, Riemann solver approaches are only benchmarked against shocks, where they perform, by construction, very well. However, such approaches can still be way too dissipative to accurately model the growth of weakly triggered fluid instabilities.
It has also turned out that the still frequently used cubic spline kernel is not a great choice after all, but substantially better kernels are readily available (Wendland, 1995; Cabezon et al., 2008; Dehnen & Aly, 2012; Rosswog, 2015a) and replacing the kernel function only requires small changes to existing codes. It has further been realized that much more accurate gradient estimates (Garcia-Senz et al., 2012; Cabezon et al., 2012; Rosswog, 2015a) than the standard kernel gradients can be achieved at a moderate additional cost and actually even without sacrificing the highly valued kernel anti-symmetry, $\nabla_{a}W_{ab}=-\nabla_{b}W_{ab}$, which is crucial for good numerical conservation.
Several of the above improvements have borrowed techniques that are traditionally used in Finite Volume schemes such as reconstruction, slope-limiting or Riemann solvers. There is however, also a class of methods that tries to formulate the ideal gas dynamics equations from the beginning as Finite Volume methods, though for particles rather than for meshes, see e.g. Ben Moussa et al. (1999); Vila (1999); Hietel et al. (2000); Junk (2003); Gaburov & Nitadori (2011); Hopkins (2015). This also comes with a broad variety of names for relatively similar methods, some of which are ”proper SPH”, some are ”SPH with elements from Finite Volume methods” and yet others are proper Finite Volume discretizations of the conservation laws written for stationary particles (Eulerian) or for particles that either move with the fluid velocity (Lagrangian) or with any other velocity (Adaptive Lagrangian Eulerian or ALE). Therefore the boundaries between different methods are blurring and it becomes a matter of semantics or taste how to call a given method. In summary, many of the issues that SPH has been criticized for, can be considered as essentially solved and modern SPH formulations can accurately solve very challenging gas dynamics problems [^1] For the publication Rosswog (2020b) the author collected in the computational astrophysics community a large set of test problems ”that SPH cannot do”, but please see the results in this paper., see for example (Frontiere et al., 2017; Rosswog, 2020b; Sandnes et al., 2024).
Figure 1: Illustration of one of SPH’s most salient features: the natural treatment of vacuum. The plot shows how two stars have been ripped apart by a black hole (lurking at the coordinate origin) into spaghetti-like, thin gas streams that are held together by the gas’ self-gravity. The inset shows the initial conditions of the initially spherical stars. The original stars in the inset only cover a fraction of $10^{-9}$ of the volume that is shown at late times (t= 371 h). Simulation performed with the MAGMA2 code, see Rosswog (2020).
One issue that has been strongly improved by the above measures, but is usually still not exactly enforced in standard SPH approaches, is the lack of zero order consistency. In simple words, the SPH approximation makes use of weighted kernel sums over neighbouring particles, but the weights in these sums do not add up exactly to unity, see the discussion in section  2.4. Whether this is an issue in practical applications or not depends on the used kernels/neighbour numbers and the considered test problem, but it is definitely desirably to avoid the issue in the first place. Otherwise, not even constant functions are reproduced exactly.
In this paper, we formulate and explore a new approach, in which we start from one of the commonly used SPH discretizations, but a) we introduce Roe’s approximate solver, b) we identify the dissipative terms in the Riemann solver and in these terms we use slope-limited reconstruction and, finally, c) we use linearly reproducing kernels so that constant and linear functions are reproduced to machine precision. We carefully scrutinize our implementation of the reproducing kernels and we put the new method to the test with a number of challenging benchmark tests, ranging from shocks over fluid instabilities to vorticity creating shocks à la Schulz-Rinne (1993). All of these tests are performed in three spatial dimensions. In Sec. 2 we discuss all the involved elements of our approach, in Sec. 3 we present and discuss our benchmark tests before we summarize and conclude in Sec. 4.
2 Methodology
2.1 Particle hydrodynamics formulation with a Riemann solver
Our starting point is a common set of Smoothed Particle Hydrodynamics (SPH) equations (Parshikov & Medin, 2002; Liu & Liu, 2003; Meng et al., 2021)
$\rho_{a}$ $=$ $\sum_{b}m_{b}\bar{W}_{ab}$ (1)
$\frac{d\vec{v}_{a}}{dt}$ $=$ $-\sum_{b}m_{b}\frac{P_{a}+P_{b}}{\rho_{a}\rho_{b}}\nabla_{a}\bar{W}_{ab}$ (2)
$\frac{du_{a}}{dt}$ $=$ $\frac{1}{2}\sum_{b}m_{b}\frac{P_{a}+P_{b}}{\rho_{a}\rho_{b}}\vec{v}_{ab}\cdot\nabla_{a}\bar{W}_{ab}.$ (3)
As usual, one has the choice between either solving the continuity equation directly, or to keep the particle masses, $m_{b}$, fixed and obtain the density $\rho_{a}$ via a kernel-weighted sum over neighbours, see Eq. (1), which is our choice here. The continuity equation approach has been found advantageous in some recent studies (Meng et al., 2021; Sandnes et al., 2024), but we choose here the summation approach since it is guaranteed to deliver a strictly positive density for the positive definite kernels $W$ that we are using. The kernel support size is set by the smoothing length $h$, $P$ denotes the gas pressure, $u$ is the specific internal energy and $\vec{v}_{ab}=\vec{v}_{a}-\vec{v}_{b}$. Often, the density estimation is performed ”as locally as possible” in the sense that the kernel in Eq. (1) is evaluated with the smoothing length of particle $a$, $h_{a}$. To be consistent with our later choices, see Sec. 2.4, we choose here
$\bar{W}_{ab}=\frac{W(|\vec{r}_{a}-\vec{r}_{b}|/h_{a})+W(|\vec{r}_{a}-\vec{r}_{b}|/h_{b})}{2}$ (4)
and
$\nabla\bar{W}_{ab}=\frac{\nabla W(|\vec{r}_{a}-\vec{r}_{b}|/h_{a})+\nabla W(|\vec{r}_{a}-\vec{r}_{b}|/h_{b})}{2}.$ (5)
In Eqs (2) and (3), the index $a$ at the nabla operator indicates that it is evaluated with respect to particle $a$ (and not $b$).
While we have written the above equations in a commonly used way, it is worth keeping in mind that the kernel gradient can explicitly written (for some smoothing length $h$) as
$\nabla_{a}W_{ab}(h)=\frac{\partial W(q_{ab})}{\partial q_{ab}}\frac{\partial q_{ab}}{\partial\vec{r}_{a}}=\frac{\partial W(q_{ab})}{\partial q_{ab}}\frac{\hat{e}_{ab}}{h},$ (6)
where $q_{ab}=|\vec{r}_{a}-\vec{r}_{b}|/h$ and $\hat{e}_{ab}=(\vec{r}_{a}-\vec{r}_{b})/|\vec{r}_{a}-\vec{r}_{b}|$ is the unit vector pointing from particle $b$ to particle $a$. With this, Eq.(3) can also be written as
$\frac{du_{a}}{dt}$ $=$ $\frac{1}{2}\sum_{b}m_{b}\frac{P_{a}+P_{b}}{\rho_{a}\rho_{b}}\left(\vec{v}_{a}\cdot\hat{e}_{ab}-\vec{v}_{a}\cdot\hat{e}_{ab}\right)\frac{1}{h_{ab}}\frac{\partial W}{\partial q_{ab}}$ (7)
$=$ $\frac{1}{2}\sum_{b}m_{b}\frac{P_{a}+P_{b}}{h_{ab}\;\rho_{a}\rho_{b}}\frac{\partial W}{\partial q_{ab}}\left(\tilde{v}_{a}-\tilde{v}_{b}\right)$ (8)
where the velocities projected on the connection line have been abbreviated with a tilde.
So far, the equations are entirely non-dissipative and in order to handle shocks they need to be enhanced by mechanisms that produce sufficient (but not too much) dissipation to obtain non-oscillatory solutions in shocks. As outlined in the introduction, a commonly approach is to use artificial viscosity which has seen many improvements in recent times (Cullen & Dehnen, 2010; Rosswog, 2015a; Frontiere et al., 2017; Rosswog, 2020a, b; Rosswog & Diener, 2021; Rosswog et al., 2022, 2023; Sandnes et al., 2024). These new approaches have turned out to be both robust and accurate and to overall produce very little unwanted dissipation. In our understanding, such modern artificial dissipation approaches are on par with approximate Riemann solvers.
Figure 2: To add dissipation, Riemann problems are solved for each particle and its neighbour particles.
Despite these successes, we want to explore here dissipation that is provided by a Riemann solver. The major idea is to solve (approximately) a one dimensional Riemann problem at the midpoint between each pair of interacting particles $a$ and $b$, as sketched in Fig. 2. In particular, we replace averages of particle values by the solution of a Riemann problem
$\frac{P_{a}+P_{b}}{2}\rightarrow P_{ab}^{\ast}\quad{\rm and}\quad\frac{\tilde{v}_{a}+\tilde{v}_{b}}{2}\rightarrow v_{ab}^{\ast},$ (9)
where the $\ast$ labels the contact discontinuity state in a Riemann problem, see e.g. Toro (1999), and $ab$-index indicates that it is the solution between state $a$ and $b$ (not to be confused with $\vec{v}_{ab}=\vec{v}_{a}-\vec{v}_{b}$). The tilde denotes velocities projected onto the line connecting two particles
$\tilde{v}_{a}=\vec{v}_{a}\cdot\hat{e}_{ab}\quad{\rm and}\quad\tilde{v}_{b}=\vec{v}_{b}\cdot\hat{e}_{ab}.$ (10)
With the substitutions Eq. (9) we have
$P_{a}+P_{b}=2P_{ab}^{\ast}\quad{\rm and}\quad\tilde{v}_{a}-\tilde{v}_{b}=2(\tilde{v}_{a}-v_{ab}^{\ast})$ (11)
so that the equations, now with dissipation, read
$\rho_{a}$ $=$ $\sum_{b}m_{b}\bar{W}_{ab}$ (12)
$\frac{d\vec{v}_{a}}{dt}$ $=$ $-\frac{2}{\rho_{a}}\sum_{b}V_{b}P_{ab}^{\ast}\nabla_{a}\bar{W}_{ab}$ (13)
$\frac{du_{a}}{dt}$ $=$ $\frac{2}{\rho_{a}}\sum_{b}V_{b}P_{ab}^{\ast}(\vec{v}_{a}-\vec{v}^{\ast}_{ab})\cdot\nabla_{a}\bar{W}_{ab},$ (14)
where we have used $V_{b}=m_{b}/\rho_{b}$ and $\vec{v}^{\ast}_{ab}=v_{ab}^{\ast}\hat{e}_{ab}$.
2.2 Dissipation in the Roe solver
We use Roe’s approximate Riemann solver (Roe, 1986) for the star state:
$v_{ab}^{\ast}$ $=$ $\frac{1}{2}\left((\vec{v}_{a}+\vec{v}_{b})\cdot\hat{e}_{ab}+\frac{P_{b}-P_{a}}{C_{\rm RL}}\right)$ (15)
$P_{ab}^{\ast}$ $=$ $\frac{1}{2}\left(P_{a}+P_{b}+C_{\rm RL}\frac{(\vec{v}_{b}-\vec{v}_{a})\cdot\hat{e}_{ab}}{C_{\rm RL}}\right),$ (16)
where the Roe-averaged Lagrangian sound speed is
$C_{\rm RL}=\frac{c_{s,a}\rho_{a}\sqrt{\rho_{a}}+c_{s,b}\rho_{b}\sqrt{\rho_{b}}}{\sqrt{\rho_{a}}+\sqrt{\rho_{b}}}.$ (17)
If, for a moment, we ignore the terms that involve the pressure and velocity differences in Eqs.(15) and (16), we have
$v_{ab}^{\ast}\approx\frac{\vec{v}_{a}+\vec{v}_{b}}{2}\cdot\hat{e}_{ab}\quad{\rm and}\quad P_{ab}^{\ast}\approx\frac{P_{a}+P_{b}}{2}.$ (18)
By inserting these expressions into Eqs. (13) and (14) we obviously recover the inviscid equations (2) and (3), therefore the terms involving the differences in Eqs. (15) and (16) are responsible for the dissipation! One may thus try to cure potentially excessive dissipation by modifying these terms. We approach this issue here by applying reconstructed pressure and velocity values in the dissipative terms, so that for perfectly reconstructed smooth flows we are solving the inviscid equations.
2.3 Reconstruction in the dissipative terms
Our strategy to reduce dissipation is to use in the dissipative terms values of $P$ and $\vec{v}$ that are reconstructed to the mid-point of each particle pair, $\vec{r}_{ab}^{\rm mid}=0.5(\vec{r}_{a}+\vec{r}_{b})$, see Fig. 2. So at each inter-particle midpoint one has $P$/$\vec{v}$ values that are reconstructed once from the $a$-side and once from the $b$-side. Explicitly, the linearly reconstructed velocity components read
$v_{a}^{i,\rm rec}$ $=$ $v^{i}_{a}-\frac{1}{2}\Phi(\nabla v^{i}_{a},\nabla v^{i}_{b})\cdot\vec{r}_{ab}$ (19)
$v_{b}^{i,\rm rec}$ $=$ $v^{i}_{b}+\frac{1}{2}\Phi(\nabla v^{i}_{a},\nabla v^{i}_{b})\cdot\vec{r}_{ab},$ (20)
the specific internal energy is
$u_{a}^{\rm rec}$ $=$ $u_{a}-\frac{1}{2}\Phi(\nabla u_{a},\nabla u_{b})\cdot\vec{r}_{ab}$ (21)
$u_{b}^{\rm rec}$ $=$ $u_{b}+\frac{1}{2}\Phi(\nabla u_{a},\nabla u_{b})\cdot\vec{r}_{ab}$ (22)
and the density
$\rho_{a}^{\rm rec}$ $=$ $\rho_{a}-\frac{1}{2}\Phi(\nabla\rho_{a},\nabla\rho_{b})\cdot\vec{r}_{ab}$ (23)
$\rho_{b}^{\rm rec}$ $=$ $\rho_{b}+\frac{1}{2}\Phi(\nabla\rho_{a},\nabla\rho_{b})\cdot\vec{r}_{ab}.$ (24)
With the reconstructed values of internal energy and density we can calculate the corresponding pressure values via our equation of state, $P=(\gamma-1)\rho u$.
The quantity $\Phi$ that is used in the above reconstructions is a suitable slope limiter. We experiment with commonly used slope limiters such as minmod
$\Phi_{mm}(x,y)=\frac{1}{2}\left[\rm{SGN}(x)+\rm{SGN}(y)\right]\;{\rm MIN}(|x|,|y|),$ (25)
or the vanLeer limiter (van Leer, 1977)
$\Phi_{vL}(x,y)=\begin{cases}\frac{2xy}{x+y}\quad&{\rm if}\;xy>0\\ 0&{\rm otherwise}.\end{cases}$ (26)
Other popular limiters are the vanLeer monotonized Central (vanLeerMC) (van Leer, 1977)
$\Phi_{vLMC}(x,y)=\begin{cases}\rm{SGN}(x)\;\rm{MIN}\left[\frac{|x+y|}{2},2|x|,2|y|\right]\quad&{\rm if}\;xy>0\\ 0&{\rm otherwise},\end{cases}$ (27)
and the vanAlbada limiter (van Albada et al., 1982)
$\Phi_{vA}(x,y)=\begin{cases}\frac{(x^{2}+\epsilon^{2})y+(y^{2}+\epsilon^{2})x}{x^{2}+y^{2}+2\epsilon^{2}}\quad&{\rm if}\;xy>0,\\ 0&{\rm otherwise}.\end{cases}$ (28)
The results are insensitive to the exact value of $\epsilon$, here, we use $\epsilon^{2}=10^{-6}$. All of the limiters are applied component-wise if the inputs are vectors.
Figure 3: Particle distribution to scrutinize the linear reproducing kernel method. The 3D particle distribution has been set up using a Centroidal Voronoi Tesselation, shown is a slice with a thickness of $z<6\times 10^{-3}$.
Refer to caption Refer to caption
Figure 4: Approximation of function value (left) and derivative in $x$-direction (right), the exact result is in both cases $=1$. Each time we show the SPH-approximation in black and the RPK-result in red.
Refer to caption Refer to caption
Figure 5: Logarithm of the function error (left) and of the error in the $x$-derivative approximation (right). Each time we show the SPH-approximation in black and the RPK-result in red.
2.4 Linearly reproducing kernel (RPK) approximations
One of the main criticisms of the standard SPH method is that it neither exactly reproduces constant nor linear functions. In the standard SPH-discretization the approximation of a function $f(\vec{r})$ reads
$\tilde{f}(\vec{r})=\sum_{b}\;V_{b}\;f_{b}\;W(\vec{r}-\vec{r}_{b},h(\vec{r}))=\sum_{b}f_{b}\;\Phi_{b}(\vec{r})$ (29)
where we have abbreviated $m_{b}/\rho_{b}$ as $V_{b}$ and the smoothing length $h$ should be thought of as a function of $\vec{r}$. After the second equal sign, we have abbreviated the product $V_{b}W(\vec{r}-\vec{r}_{b},h(\vec{r}))$ as weight function $\Phi_{b}(\vec{r})$. As an example, if all function values would be the same, $f_{b}=f_{0}$, then the function approximation should yield $f_{0}$, but the standard SPH approximation finds
$\tilde{f}(\vec{r})=f_{o}\sum_{b}\frac{m_{b}}{\rho_{b}}W(\vec{r}-\vec{r}_{b},h)\approx f_{0},$ (30)
which is approximately, but not exactly, equal to the desired value, because the sum is not guaranteed to yield exactly unity. More systematically, one can expand $f_{b}$ (i.e. the function $f$ at the position $b$) around $\vec{r}$, insert this into Eq. (29) to find
$\tilde{f}(\vec{r})$ $=$ $\sum_{b}\left[f(\vec{r})+(\nabla f)(\vec{r})\cdot(\vec{r_{b}}-\vec{r})+{\rm h.o.t.}\right]\;\Phi_{b}(\vec{r})$ (31)
$=$ $f(\vec{r})\sum_{b}\Phi_{b}(\vec{r})+(\nabla f)(\vec{r})\cdot\sum_{b}(\vec{r_{b}}-\vec{r})\Phi_{b}(\vec{r})+{\rm h.o.t.},$ (32)
which implies, not too surprisingly, that for a good approximation, $\tilde{f}(\vec{r})\approx f(\vec{r})$, the lowest order consistency relations
$\sum_{b}\Phi_{b}(\vec{r})=1\quad{\rm and}\quad\sum_{b}(\vec{r}_{b}-\vec{r})^{i}\;\Phi_{b}(\vec{r})=0,$ (33)
should be fulfilled. Standard SPH usually fulfills this well in initial conditions, but it does not enforce these conditions during the evolution. The conditions are well fulfilled when good kernels (e.g. of the Wendland family (Wendland, 1995)) with large neighbour numbers are used, but this, of course, comes at the price of expensive sums over neighbouring particles.
The consistency relations Eqs. (33), however, can also be enforced by construction, e.g. in the so-called reproducing kernel method (Liu et al., 1995). One can, for example, enhance the kernel functions with additional parameters $A$ and $B^{i}$,
$\mathcal{W}_{ab}(\vec{r}_{ab})=A_{a}\left[1+B_{a}^{i}\;(\vec{r}_{ab})^{i}\right]\bar{W}_{ab},$ (34)
where $\bar{W}_{ab}$ is given by Eq. (4), and then determine the four unknown numbers $A$ and $B^{i}$ by enforcing the four discrete consistency relations so that
$\sum_{b}V_{b}\;\mathcal{W}_{ab}=1\quad{\rm and}\quad\sum_{b}V_{b}\;(\vec{r}_{ab})^{i}\;\mathcal{W}_{ab}=0$ (35)
is fulfilled at every particle position $\vec{r}_{a}$. Since these kernels $\mathcal{W}_{ab}$ reproduce by construction constant and linear functions exactly (i.e. to floating point accuracy), they are usually referred to as (linearly) reproducing kernels (RPKs). Reconstructing kernels for higher order polynomials could be designed in a similar fashion, but at the price of lengthy and computationally expensive to evaluate expressions. We therefore restrict ourselves here to linear order. Since the $B^{i}$ are in general non-zero, the kernels $\mathcal{W}_{ab}$ are not radial as standard SPH kernels, and therefore angular momentum conservation is difficult to guarantee. In SPH angular momentum conservation is usually a consequence of the interparticle forces pointing along the connecting vectors between a particle $a$ and a particle $b$, $\hat{e}_{ab}$, and the kernel being anti-symmetric, $\nabla_{a}W_{ab}=-\nabla_{b}W_{ab}$, see Sec. 2.4 in Rosswog (2009) for a detailed discussion. It is, however, still possible to write a set of equations that conserves energy and momentum (and mass if density summation is used), but not necessarily angular momentum. In practice, however, this does not seem to be a major concern, Frontiere et al. (2017) find with their artificial viscosity based RPK-SPH approach in typical tests violations of exact angular momentum conservation on the sub-percent level.
The gradient of $\mathcal{W}_{ab}$ can be calculated in a straight-forward way as
$(\nabla_{a})^{k}\mathcal{W}_{ab}$ $=$ $A_{a}\;B_{a}^{k}\;\bar{W}_{ab}+A_{a}\left(1+B_{a}^{i}(\vec{r}_{ab})^{i}\right)\nabla_{a}^{k}\bar{W}_{ab}+$ (36)
$\left(1+B_{a}^{i}(\vec{r}_{ab})^{i}\right)\bar{W}_{ab}(\partial_{k}A)_{a}+A_{a}\;(\vec{r}_{ab})^{i}\;(\partial_{k}B)_{a}^{i}\;\bar{W}_{ab}$
Now taking the nabla-operator with respect to particle $b$ one finds
$(\nabla_{b})^{k}\mathcal{W}_{ba}=$ $A_{b}\;B_{b}^{k}\;\bar{W}_{ba}+A_{b}\left(1+B_{b}^{i}(\vec{r}_{ba})^{i}\right)\nabla_{b}^{k}\bar{W}_{ba}+$
$\left(1+B_{b}^{i}(\vec{r}_{ba})^{i}\right)\bar{W}_{ba}(\partial_{k}A)_{b}+A_{b}\;(\vec{r}_{ba})^{i}\;(\partial_{k}B)_{b}^{i}\;\bar{W}_{ba}$
$=$ $A_{b}\;B_{b}^{k}\;\bar{W}_{ab}+A_{b}\left(1+B_{b}^{i}(\vec{r}_{ab})^{i}\right)\nabla_{a}^{k}\bar{W}_{ab}+$ (37)
$\left(1-B_{b}^{i}(\vec{r}_{ab})^{i}\right)\bar{W}_{ab}(\partial_{k}A)_{b}-A_{b}\;(\vec{r}_{ab})^{i}\;(\partial_{k}B)_{b}^{i}\;\bar{W}_{ab},$
where we have used $\bar{W}_{ab}=\bar{W}_{ba}$, $\vec{r}_{ba}=-\vec{r}_{ab}$ and $\nabla_{b}^{k}\bar{W}_{ba}=-\nabla_{a}^{k}\bar{W}_{ab}$. The parameters $A$ and $B^{i}$ and their derivatives, which are needed for $\partial_{k}\mathcal{W}_{ab}$, can then be calculated by straight-forward algebra. We first define discrete moments at position $\vec{r}_{a}$ as
$\left(M_{0}\right)_{a}$$\equiv$ $\sum_{b}V_{b}\;\bar{W}_{ab}$ (38)
$\left(M_{1}^{i}\right)_{a}$$\equiv$ $\sum_{b}V_{b}\;(\vec{r}_{ab})^{i}\bar{W}_{ab}$ (39)
$\left(M_{2}^{ij}\right)_{a}$$\equiv$ $\sum_{b}V_{b}\;(\vec{r}_{ab})^{i}\;(\vec{r}_{ab})^{j}\;\bar{W}_{ab}$ (40)
and their derivatives read
$\left(\partial_{k}M_{0}\right)_{a}$$\equiv$ $\sum_{b}V_{b}\;\nabla_{a}^{k}\bar{W}_{ab}$ (41)
$\left(\partial_{k}M_{1}^{i}\right)_{a}$$\equiv$ $\sum_{b}V_{b}\left[(\vec{r}_{ab})^{i}(\nabla_{a}^{k}\bar{W}_{ab})+\delta^{ki}\bar{W}_{ab}\right]$ (42)
$\left(\partial_{k}M_{2}^{ij}\right)_{a}$$\equiv$ $\sum_{b}V_{b}\;[(\vec{r}_{ab})^{i}\;(\vec{r}_{ab})^{j}\;(\nabla_{a}^{k}\bar{W}_{ab})+(\vec{r}_{ab})^{i}\;\delta^{jk}\;\bar{W}_{ab}+$ (43)
$(\vec{r}_{ab})^{j}\;\delta^{ik}\;\bar{W}_{ab}].$
With the moments and their derivatives at hand, one can calculate the kernel parameters
$A_{a}$ $=$ $\frac{1}{\left(M_{0}\right)_{a}-(M_{2}^{ij})^{-1}_{a}\;\left(M_{1}^{i}\right)_{a}\;(M_{1}^{j})_{a}}$ (44)
$B_{a}^{i}$ $=$ $-(M_{2}^{ij})_{a}^{-1}\;(M_{1}^{j})_{a}$ (45)
and their somewhat lengthy, but straight forward calculable derivatives
$\partial_{k}A_{a}=$ $-A^{2}_{a}[(\partial_{k}M_{0})_{a}-2(M_{2}^{ij})_{a}^{-1}(M_{1}^{j})_{a}\left(\partial_{k}M_{1}^{i}\right)_{a}$ (46)
$+(M_{2}^{il})_{a}^{-1}(\partial_{k}M_{2}^{lm})_{a}(M_{2}^{mj})_{a}^{-1}(M_{1}^{j})_{a}(M_{1}^{i})_{a}]$
and
$\partial_{k}B_{a}^{i}=-(M_{2}^{ij})_{a}^{-1}\;(\partial_{k}M_{1}^{j})_{a}+(M_{2}^{il})_{a}^{-1}\;(\partial_{k}M_{2}^{lm})_{a}(M_{2}^{mj})_{a}^{-1}(M_{1}^{j})_{a}.$ (47)
With the linearly reproducing kernels at hand, we can now approximate a function $F$ via
$\tilde{F}(\vec{r}_{a})=\sum_{b}V_{b}\;F_{b}\;\mathcal{W}_{ab}$ (48)
and its derivative
$\partial_{k}\tilde{F}(\vec{r}_{a})=\sum_{b}V_{b}\;F_{b}\;\partial_{k}\mathcal{W}_{ab}.$ (49)
While these expressions look very similar to standard SPH approximations, they exactly reproduce linear functions on a discrete level, which the SPH equations do not.
2.4.1 Scrutinizing the implementation
To test our implementation of the linear reproducing kernels, we start from a relatively regular, but not exactly uniform particle distribution, sometimes referred to as ”a glass”. Our initial particle configuration, see Fig. 3, has been produced via a Centroidal Voronoi Tesselation (CVT; Du, Faber & Gunzburger 1999) in the computational volume [0.5,0.5] $\times$ [0.5,0.5] $\times$ [0.5,0.5], and it is further regularized by additional sweeps according to the ”artificial pressure method” (APM), as described in the MAGMA2 paper, see especially Eq.(40). We perform the regularisation sweeps for the inner regions while the outer regions remain as boundary particles on the original CVT setup. The resulting particle distribution is most regular in the centre, and approaching the somewhat rougher original distribution as one moves away from the centre. We now assign functions $f$ to each particle, a) once a constant value $f^{(1)}=1$ and b) the other time we assign a linear function that increases in $x$-direction $f^{(2)}=x$ with a slope of unity.
Since by construction, the consistency relations, Eq. (35), are fulfilled at the particle positions (i.e. not everywhere in space), we select every 100th particle in the inner regions (so that the absolute value of each coordinate is $<0.4$) and at these particles, labelled $a$, we calculate the standard SPH approximations
$f^{\rm SPH}_{a}=\sum_{b}V_{b}f_{b}\bar{W}_{ab}\quad{\rm and}\quad\left(\nabla f\right)^{\rm SPH}_{a}=\sum_{b}V_{b}f_{b}\nabla\bar{W}_{ab}$ (50)
as well as the reproducing kernel approximations
$f^{\rm RPK}_{a}=\sum_{b}V_{b}f_{b}\mathcal{W}_{ab}\quad{\rm and}\quad\left(\nabla f\right)^{\rm RPK}_{a}=\sum_{b}V_{b}f_{b}\nabla\mathcal{W}_{ab}$ (51)
and compare against the theoretical results of unity for both the function value and the derivative in $x$-direction. For the SPH-type kernel $W$ that enters Eq. (4) we choose, here and in the rest of the paper, a member of the family of harmonic-like kernels (Cabezon et al. (2008))
$W(x)=W^{\rm H}_{8}(x)=\frac{\sigma_{8}}{h^{3}}\left\{\begin{array}[]{llr}1,&\;x=0,\\[11.38109pt] \left(\frac{\sin\left[\frac{\pi}{2}x\right]}{\frac{\pi}{2}x}\right)^{8}&\;0<x<2,\\[11.38109pt] 0,&\;{\rm else},\end{array}\right.$ (55)
[…]
In all fairness, it is worth stating that the SPH-approximation results for this particle distribution look rather poor, but the question is whether such noisy particle distributions actually occur in an SPH simulation in the first place. Much recent work has been invested in designing simulations techniques so that this does not happen, for example by using Wendland kernels with large neighbour number (typically several hundreds), see e.g. Rosswog 2015a,b. This typically produces very regular particle distributions where the error is much smaller than the ones shown here. The errors, however, will never be smaller than what one finds for a regular lattice ($\sim 10^{-5}$). So even for close-to-perfect particle distributions, the RPK-approximations will be more than nine orders of magnitude more accurate than the standard SPH-approach.
Figure 6: Spherical blast wave problem1: density, velocity and pressure, vanAlbada limiter.
Refer to caption Refer to caption Refer to caption
Figure 7: Spherical blast wave problem1: density, velocity and pressure with a strip around the x axis. For this simulation $200^{3}$ particles and the vanAlbada limiter were used, the reference solution (red line) was obtained by the Eulerian-weighted average flux method ($400^{3}$ grid cells, Toro 1999).
2.5 Final equation set
Figure 8: Spherical blast wave problem2: density, velocity and pressure, vanAlbada limiter.
Refer to caption Refer to caption Refer to caption
Figure 9: Spherical blast wave problem2: density, velocity and pressure with a strip around the x axis. For this simulation $200^{3}$ particles and the vanAlbada limiter were used, the reference solution (red line) was obtained by the Eulerian-weighted average flux method ($400^{3}$ grid cells, Toro 1999).
Figure 10: Spherical blast wave problem 2 (detail): shown is the comparison of the densities for different slope limiters.
One desirable quantity for numerical conservation is the anti-symmetry of the kernel gradient
$\nabla_{a}W_{ab}(h_{ab})=-\nabla_{b}W_{ab}(h_{ab}),$ (56)
which in typical SPH equations guarantees in a straight forward way the numerical conservation of energy and momentum, since in the time derivative of their total values all terms cancel exactly, see for example Sec. 2.4 in Rosswog (2009). The usual standard kernel gradients point in the direction of the line connecting two particles, see Eq. (6), therefore together with the kernel gradient anti-symmetry also the exact conservation of angular momentum can be guaranteed. Due to the presence of the vector $B^{i}$ in Eq. (34) the kernels $\mathcal{W}$ are no longer guaranteed to be radial and therefore angular momentum conservation cannot be guaranteed in the same way as in standard SPH. The standard kernel gradient as it enters Eqs.(13) and (14) can also be written as
$\nabla_{a}\bar{W}_{ab}=\frac{\nabla_{a}W_{ab}(h_{a})-\nabla_{b}W_{ba}(h_{b})}{2},$ (57)
where we have made use of the anti-symmetry of the kernel gradients. We now replace the gradients of the kernels $W$ by the gradients of the much more accurate gradients of the kernels $\mathcal{W}$, see Eqs.(36) and (37), or, in other words, we replace $\nabla_{a}\bar{W}_{ab}$ by
$(\nabla\mathcal{W})_{ab}\equiv\frac{1}{2}\left[\nabla_{a}\mathcal{W}_{ab}-\nabla_{b}\mathcal{W}_{ba}\right],$ (58)
so that our final equation set reads
$\rho_{a}$ $=$ $\sum_{b}m_{b}\bar{W}_{ab}$ (59)
$\frac{d\vec{v}_{a}}{dt}$ $=$ $-\frac{2}{\rho_{a}}\sum_{b}V_{b}P_{ab}^{\ast}(\nabla\mathcal{W})_{ab}$ (60)
$\frac{du_{a}}{dt}$ $=$ $\frac{2}{\rho_{a}}\sum_{b}V_{b}P_{ab}^{\ast}(\vec{v}_{a}-\vec{v}^{\ast}_{ab})\cdot(\nabla\mathcal{W})_{ab}.$ (61)
We also use the gradients $(\nabla\mathcal{W})_{ab}$ in the reconstruction process described in Sec. 2.3. As mentioned before, our equation set does not manifestly conserve angular momentum as standard SPH does, but, at least in the below test cases, we find excellent results with our approach.
3 Results
We will explore the performance of the equations (59) to (61) in a set of challenging benchmarks involving shocks, instabilities and complex shocks with vorticity creation à la Schulz-Rinne (1993). We will each time also explore the performance of the following slope limiters (in order decreasing dissipation): minmod, vanAlbada and vanLeerMC.
3.1 Spherical blast wave 1
We start with a three-dimensional shock-tube-type problem. We choose the same parameters as Toro (1999) (apart from a shift of the origin): the computational domain is $[-1,1]^{3}$ and the initial conditions are chosen as:
$(\rho,\vec{v},P)=\left\{\begin{array}[]{l}(1.000,0,0,0,1.0)\quad{\rm for\;\;r<0.5}\\ (0.125,0,0,0,0.1)\quad{\rm else.}\end{array}\right.$ (62)
The solution exhibits a spherical shock wave, a spherical contact surface traveling in the same direction and a spherical rarefaction wave traveling towards the origin. As initial conditions we simply placed $200^{3}$ particles on a cubic lattice within $[-1,1]^{3}$, together with surounding, ”frozen” particles as boundary condition.
We show in Fig. 6 the values of density, velocity and pressure for the vanAlbada limiter. Despite the initial setup on a cubic lattice, the results are practically perfectly spherically symmetric. The results for the different limiters are merely identical, only on very close inspection, one can find a reminiscence of the grid structure in the case of the vanLeerMC limiter. In Fig. 7 we show our particle results in a thin strip around the $x-$axis ($|y|<0.018,|z|<0.018$) compared with a $400^{3}$ grid cell calculation with the Eulerian-weighted average flux method (Toro, 1999). Overall, we find excellent agreement. This figure can be compared to Fig.  13 of Rosswog (2020b) where the same test was performed at the same resolution, but with the MAGMA2 code using matrix inversion gradients and an artificial viscosity prescription that also uses slope limited reconstruction. For more details on the latter method see the ”Matrix Inversion method I” in Rosswog (2020b). While both approaches show excellent agreement with the reference solution, the RPK-approach of this paper yields a sharper resolved contact discontinuity.
3.2 Spherical blast wave 2
As a second spherical blast wave problem (Toro, 1999) we start from
$(\rho,\vec{v},P)=\left\{\begin{array}[]{l}(1.0,0,0,0,2.0)\quad{\rm for\;\;r<0.5}\\ (1.0,0,0,0,1.0)\quad{\rm else}\end{array}\right.$ (63)
and place again $200^{3}$ particles in the same straight forward way as in the first blast wave problem.
We show the numerical solution of this test in Fig. 8. Again, deviations from sphericity are minute and the agreement with the reference solution (Eulerian-weighted average flux method with $400^{3}$ grid cells (Toro, 1999)), see Fig. 9, is excellent. In Fig. 10 we show a zoom into the density distribution to illustrate the effect of the different slope limiters. Overall the agreement is very good, but as expected, the minmod limiter is the most diffusive one, the vanLeerMC limiter captures best the edges of the rarefaction wave, but at the price of a small overshoot at the shock front. The results can again be compared to the ”Matrix Inversion method I”, see Figs.14 and 15 in Rosswog (2020b): overall the results are very similar, but the RPK-hydrodynamics does not show a density overshoot in the central region which the older result does.
Figure 11: Density evolution ($XY$-pane) in a Sedov blast wave test with the vanAlbada slope limiter.
Figure 12: Impact of the slope limiter on a Sedov blast wave: the result for minmod is shown in the left panel, for vanAlbada in the middle and vanLeerMC in the right panel.
3.3 Sedov blast
The Sedov-Taylor explosion test, a strong, initially point-like blast expanding into a low density environment, has an analytic self-similarity solution (Sedov, 1959; Taylor, 1950). For an explosion energy $E$ and a density of the ambient medium $\rho$, the blast wave propagates after a time $t$ to the radius $r(t)=\beta(Et^{2}/\rho)^{1/5}$, where $\beta$ depends on the adiabatic exponent of the gas ($\approx 1.15$ in 3D for the $\Gamma=5/3$ that we are using). In the strong explosion limit, the density jumps at the shock front by a factor of $\rho_{2}/\rho_{1}=(\Gamma+1)/(\Gamma-1)=4$, where the numerical value refers to our chosen value of $\Gamma$. Behind the shock the density drops quickly and finally vanishes at the centre of the explosion.
To set up the test numerically, we distribute $256^{3}$ SPH particles according to a Centroidal Voronoi Tesselations (CVT) (Du et al., 1999) in the computational volume [-0.5,0.5]$\times$[-0.5,0.5]$\times$[-0.5,0.5]. While this produces already reasonably good initial conditions, they can be further improved by additional sweeps according our ”Artificial Pressure Method” (APM) (Rosswog, 2020b). The main idea of the APM is to push each particle in a position where it minimizes its density error. Practically, this is achieved via the following steps: i) start from a guessed particle distribution, ii) measure at every particle the density via Eq.(1), iii) calculate the error in the density compared to a desired density profile, iv) from this error construct an artificial pressure and v) use this artificial pressure in a momentum-type equation to push the particle in a direction where the error decreases. After many such iteration steps the particles end up in locations where they optimally approximate the desired density profile. Here we need a uniform density distribution, but the method also works for rather complicated density profiles, see, for example, Fig. 3 in Rosswog (2020b). Even if the differences in the particle distributions are hard to see by eye, they still improve the outcome of the Sedov test. We therefore use here 500 of such APM-sweeps. Once the particle distribution is settled, we assign masses so that their density is $\rho=1$. This is done in an iterative way where we first assign a guess value for the masses, then measure the resulting density via Eq. (1) and subsequently correct the particle masses. The iteration is stopped once the density agrees everywhere to better than 0.5% with the desired value. The energy $E=1$ is spread across a very small initial radius $R$ and it is entirely distributed as internal energy, the specific internal energy $u$ of the particles outside of $R$ is entirely negligible ($10^{-10}$ of the central $u$). For the initial radius $R$ we choose twice the interaction radius of the innermost SPH particle. Boundaries play no role in this test as long as the blast does not interact with them. We therefore place ”frozen” particles around the computational volume as boundary particles.
In Fig. 11 we show the density evolution (cut through $XY$-pane) of the case with the vanAlbada limiter. No deviation from spherical symmetry is visible and the numerical solution for the shock from agrees very well with the exact solution (the hard-to-see black line that separates the background density (cyan) from the shocked matter). The density as a function of radius is shown in Fig. 12 where each panel shows the result for one limiter. As in the tests before, and some of the later tests, the vanAlbada limiter performs best. While here minmod performs similar, the vanLeerMC limiter does not provide enough dissipation and leads to a more noisy post-shock region and an overshoot at the shock front. Again, these results can be compared against the MAGMA2 results and -taken at face value- show slightly better agreement with the exact solution. This may, however, at least in part be due to the fact that Rosswog (2020b) we used a Wendland kernel with 300 neighbors while here we use only 220 neighor particles.
Figure 13: Kelvin-Helmholtz test with vanAlbada limiter for different resolutions at t=2.3 ($\approx 2.2\tau_{\rm KH}$).
3.4 Banded Kelvin-Helmholtz instability
Kelvin-Helmholtz (KH) shear instabilities occur in a broad range of environments, e.g. in mixing processes in novae (Casanova et al., 2011), the amplification of magnetic fields in neutron star mergers (Price & Rosswog, 2006; Giacomazzo et al., 2015; Kiuchi et al., 2015) or planetary atmospheres (Johnson et al., 2014), to name just a few examples. Traditional versions of SPH have been shown to struggle with weakly triggered KH-instabilities (Agertz et al., 2007; McNally et al., 2012), though a number of recent studies with more sophisticated numerical methods yielded very good results (Frontiere et al., 2017; Rosswog, 2020b; Sandnes et al., 2024) also for this test. We focus here on a test setup in which traditional SPH has been shown to fail, even at rather high resolution in 2D, see McNally et al. (2012). We follow the latter paper (similar setups were used in Frontiere et al. (2017) and Sandnes et al. (2024)), but we use the full 3D code and set up the ”2D” test as a thin 3D slice with $N\times N\times 20$ particles (referred to as "$N^{2}"$), for simplicity the particles are initially placed on a cubic lattice. Periodic boundary conditions are obtained by placing appropriate particle copies outside of the ”core” volume. The test is initialized as:
[…]
where $\rho_{1}=1$, $\rho_{2}=2$, $\rho_{m}=(\rho_{1}-\rho_{2})/2$ and $\Delta=0.025$. The velocity is set up as
[…]
with $v_{1}$= 0.5, $v_{2}=-0.5$, $v_{m}=(v_{1}-v_{2})/2$ and a small velocity perturbation in $y$-direction is introduced as $v_{y}=0.01\sin(2\pi x/\lambda)$ with the perturbation wave length $\lambda=0.5$. In the linear regime, a Kelvin-Helmholtz instability grows in the incompressible limit on a characteristic time scale of
$\tau_{\rm KH}=\frac{(\rho_{1}+\rho_{2})\lambda}{\sqrt{\rho_{1}\rho_{2}}|v_{1}-v_{2}|},$ (66)
with $\tau_{\rm KH}\approx 1.06$ for the chosen parameters. For our tests we choose again a polytropic equation of state with exponent $\Gamma=5/3$.
Figure 14: Kelvin-Helmholtz test ($512^{2}$) with vanAlbada limiter.
Figure 15: Kelvin-Helmholtz test ($512^{2}$) at t= 3 with different slope limiters.
Refer to caption Refer to caption
Figure 16: Growth of the Kelvin-Helmholtz instability as a function of resolution (vanLeerMC limiter, left) and for a fixed, low resolution ($128^{2}$) for different slope limiters.
We show in Fig. 13 the density at $t=2.3$ ($\approx 2.2\;\tau_{\rm KH}$) for different resolutions. Even at the lowest resolution of $128^{2}$ particles we see healthy growth that is already very similar to the better resolved cases. The evolution of the $512^{2}$ case is shown in Fig. 14. In Fig. 15 we show the density at t=3 for the different limiters. As expected, the minmod limiter is the most dissipative one and the vanLeerMC limiter is least dissipative. Similar to the Sedov test case, however, one may wonder whether the vanLeerMC limiter allows for enough dissipation, since the results look somewhat noisy.
For comparison, traditional SPH implementations struggle with this only weakly triggered instability ($v_{y}=0.01$), see, for example, Fig. 9 of McNally et al. (2012), where even at a resolution 5122 particles the instability either hardly grows (for the cubic spline kernel, label ”Ne512”) or much too slowly (for the quintic spline kernel, label ”No512”). In Fig. 16, left panel, we show the mode growth (calculated exactly as in McNally et al. (2012)) for three different resolutions, all using the vanLeerMC limiter. Their growth rates are compared to a high-resolution reference solution ($4096^{2}$ cells) obtained by the PENCIL code (Brandenburg & Dobler, 2002). Even our low-resolution case with $128^{2}$ particles is reasonably close to the reference solution. In the right panel of Fig. 16, we show the impact of the slope limiter at the example of the lowest resolution runs. As expected, the most diffusive minmod limiter grows slowest, and, once again, the vanAlbada case performs best.
Figure 17: Density in cylindrical Kelvin-Helmholtz test for the different slope limiters.
Figure 18: Rayleigh-Taylor instability test (at t=4) with vanAlbada limiter at different resolution : $128^{2}$ (left), $256^{2}$ (middle) and $512^{2}$ (right).
Figure 19: Rayleigh-Taylor instability test (at t=4.3) with $256^{2}$ particles and different slope limiters.
3.5 Cylindrical Kelvin-Helmholtz instability
As a variant of the Kelvin-Helmholtz instability, we also perform a test in cylindrical symmetry, similar to Duffell (2016). To this end we place particles from a radius of $R_{\rm BD1}=0.5$ to $R_{\rm BD2}=1.5$. For the region with $r<1$, we use
$\rho=1\quad\omega=2\quad P=4+2r^{2}$ (67)
and we use
$\rho=2\quad\omega=1\quad P=5+r^{2}$ (68)
otherwise, where $\omega$ is the angular frequency. In addition, we impose a small radial velocity perturbation to the interface given by
$v_{\rm r}=v_{0}\cos(10\;\varphi)\;e^{-\frac{(r-1)^{2}}{2\sigma_{0}^{2}}},$ (69)
where $\varphi$ is the azimuthal angle and $v_{0}=0.02$ and $\sigma_{0}=0.1$
In our initial setup we place 1.95 million particles in a uniform close-packed lattice so that the particles have the above properties. Again, we use the full 3D code to simulate a slice with 20 layers of particles in $Z$-direction, we do not allow for motion in $Z$-direction and we enforce the above velocities within $R_{\rm BD1}$ and $R_{\rm BD1}+0.05$ and $R_{\rm BD2}-0.05$ and $R_{\rm BD2}$ as boundary conditions.
We show in Fig. 17 the density at $t=2.2$, again for our three limiters. There are small artifacts at the inner and outer boundaries because of our simple treatment of the boundary conditions, but in all cases we see healthily growing Kelvin-Helmholtz instabilities with nicely winding billows. For the most diffusive limiter (minmod) we see essentially only the triggered modes growing, while in the other two cases also ”parasitic” modes have developed that have been triggered by the granularity of our closely-packed particles shearing against each other.
3.6 Rayleigh-Taylor instability
The Rayleigh-Taylor instability is a standard probe of the subsonic growth of a small perturbation. In its simplest form, a layer of density $\rho_{t}$ rests on top of a layer with density $\rho_{b}<\rho_{t}$ in a constant acceleration field, e.g. due to gravity. While the denser fluid sinks down, it develops a characteristic, ”mushroom-like” pattern. Simulations with traditional SPH implementations have shown only retarded growth or even a complete suppression of the instability (Abel, 2011; Saitoh & Makino, 2013).
We adopt again a quasi-2D setup and use the full 3D code for the evolution. We place the particles on a CL in the XY-domain $[-0.25,0.25]\times[0,1]$ and we use 20 layers of particles in the Z-direction and also place 20 layers of particles as boundaries around this core region. The properties of particles below $y=0$ are ”frozen” at the initial conditions, for the particles with $y_{a}>1$ we multiply a damping factor
$\xi=\exp\left\{-\left(\frac{y_{a}-1}{0.05}\right)^{2}\right\}$ (70)
to the temporal derivatives, so that any evolution in this upper region is strongly suppressed. We apply periodic boundaries in $x-$direction at $x=\pm 0.25$. Similar to Frontiere et al. (2017) we use $\rho_{t}=2$, $\rho_{b}=1$, a constant acceleration $\vec{g}=-0.5\hat{e}_{y}$ and
$\rho(y)=\rho_{b}+\frac{\rho_{t}-\rho_{b}}{1+\exp[-(y-y_{t})/\Delta]}$ (71)
with transition width $\Delta=0.025$ and transition coordinate $y_{t}=0.5$. We apply a small velocity perturbation to the interface
$v_{y}(x,y)=\delta v_{y,0}[1+\cos(8\pi x)][1+\cos(5\pi(y-y_{t}))]$ (72)
for $y$ in $[0.3,0.7]$ with an initial amplitude $\delta v_{y,0}=0.025$, and we use a polytropic equation of state with exponent $\Gamma=1.4$. The equilibrium pressure profile is given by
$P(y)=P_{0}-g\rho(y)[y-y_{t}]$ (73)
with $P_{0}=\rho_{t}/\Gamma$, so that the sound speed is near unity in the transition region.
We show in Fig. 18 the results for different resolutions at $t=4$ for the vanAlbada limiter. At all resolution the instability grows and reaches down to similar $Y$-values, but of course with finer detailed structures for higher resolution. In Fig. 19 we show the medium resolution case for the different limiters, again confirming the hierarchy of dissipation levels.
Figure 20: Schulz-Rinne test 1 (first row), test 2 (second row) and test 3 (third row), color-coded is the density. The first column each time refers to minmod, the second to vanAlbada and the third one to vanLeerMC.
3.7 Complex shocks with vorticity creation
Table 1: Initial data for the Schulz-Rinne-type 2D Riemann problems with vorticity creation.
SR1; contact point: $(0.3,0.3)$
variable NW NE SW SE
$\rho$ 0.5323 1.5000 0.1380 0.5323
$v_{x}$ 1.2060 0.0000 1.2060 0.0000
$v_{y}$ 0.0000 0.0000 1.2060 1.2060
$P$ 0.3000 1.5000 0.0290 0.3000
SR2; contact point: $(0.0,0.0)$
variable NW NE SW SE
$\rho$ 0.5313 1.0000 0.8000 0.5313
$v_{x}$ 0.8276 0.1000 0.1000 0.1000
$v_{y}$ 0.0000 0.0000 0.0000 0.7276
$P$ 0.4000 1.0000 0.4000 0.4000
SR3; contact point: $(0.0,0.0)$
variable NW NE SW SE
$\rho$ 1.0000 0.5313 0.8000 1.000
$v_{x}$ 0.7276 0.0000 0.0000 0.0000
$v_{y}$ 0.0000 0.0000 0.0000 0.7262
$P$ 1.0000 0.4000 1.0000 1.0000
A set of challenging 2D benchmark tests has been suggested by Schulz-Rinne (1993). The tests are constructed in such a way that four constant states meet at one corner and the initial values are chosen so that one elementary wave, either a shock, a rarefaction or a contact discontinuity appears at each interface. During the subsequent evolution complex wave patterns emerge for which no exact solutions are known. These tests are considered very challenging benchmarks for multi-dimensional hydrodynamics codes (Schulz-Rinne, 1993; Lax & Liu, 1998; Kurganov & Tadmor, 2002; Liska & Wendroff, 2003). Such tests have rarely been shown for SPH, we are only aware of the work by Puri & Ramachandran (2014) who show results for one such shock test in a study of Godunov SPH with approximate Riemann solvers, and as benchmarks for the MAGMA2 code (Rosswog, 2020b).
Here we investigate three such configurations. We simulate, as before, a 3D slice thick enough so that the midplane is unaffected by edge effects (we use 20 particle layers in Z-direction). We place particles on a cubic lattice so that 300 x 300 particles are within $[x_{c}-0.3,x_{c}+0.3]\times[y_{c}-0.3,y_{c}+0.3]$, where $(x_{c},y_{c})$ is the contact point of the quadrants, and we use a polytropic exponent $\Gamma=1.4$ in all of the tests. We refer to these Schulz-Rinne type problems as SR1 - SR3 and give their initial parameters for each quadrant in Tab. 1. These test problems correspond to configuration 3, 11 and 12 in the labelling convention of Kurganov & Tadmor (2002). We smooth the initial conditions via $\langle A\rangle_{a}=\sum_{b}\left[(m_{b}/\rho_{b})A_{b}W_{ab}\right]/\left[\sum_{b}(m_{b}/\rho_{b})W_{ab}\right]$ so that they have a smoothness that is consistent with the evolution code.
For all three tests we find very good agreement with the results published in the literature (Schulz-Rinne, 1993; Lax & Liu, 1998; Kurganov & Tadmor, 2002; Liska & Wendroff, 2003; Rosswog, 2020b). Comparing the performance of the different limiters for test SR1 (first row in Fig. 20), we see only minor differences. All cases nicely produce the shock surfaces and all show the ”mushroom-like” structure near (0.2,0.2) with only small differences. The structure around (0.3,0.3), however, does show some differences: it looks ”mushroom-like” for the least dissipative case (vanLeerMC), but not very much so for the more dissipative limiters. Also test SR2 agrees overall well with literature results, but also here the ”mushroom-like” structures are different: the curls show more windings for lower dissipation, i.e. the smallest number for minmod and the highest for vanLeerMC. Also test SR3 agrees well with literature results, here the results for the different limiters are nearly identical.
4 Summary
We have presented and explored here a new particle hydrodynamics formulation. We started from a common SPH formulation and enhanced it with Roe’s approximate Riemann solver (Roe, 1986) to solve Riemann problems between interacting particle pairs. More specifically, we replaced average values of pressures and (projected) velocities by the contact discontinuity values, $P^{\ast}$ and $v^{\ast}$. In the terms that are responsible for the dissipation, we replace the particle properties with values that are found by slope-limited reconstruction from both sides to the inter-particle midpoint. Wherever needed, we use kernels that are designed to enforce consistency relations and which, by construction, can recover constant and linear functions to machine machine precision, so-called ”reproducing kernels”. We have carefully scrutinized our corresponding implementation based on a ”glass-like” particle distribution and we compare function and gradient reproduction to the standard SPH-approximation. We find that the errors in this test are at least nine orders of magnitude lower for the RPK- than for the standard SPH-approach. These reproducing kernels are used both for the hydrodynamic gradients and in the reconstruction process.
We have put our new formulation to the test in a number of challenging benchmarks ranging from shocks over instabilities to vorticity creating shock tests designed by Schulz-Rinne (1993). We find good results in all tests, and, where comparable, we find slightly better but overall similar results to the MAGMA2 code. We have also explored three different slope limiters in the reconstruction (minmod, vanAlbada and vanLeerMC). While minmod is an overall robust choice, it leads to more dissipation than seems necessary. vanLeerMC in contrast, produces the least dissipation of the three limiters, but maybe not enough in the Sedov blast wave test where it leads to post-shock noise and an overshoot of the density. Also in our banded Kelvin-Helmholtz test, where we carefully measure the growth of the instability and compare against an extremely high-resolution calculation, we find that the vanAlbada limiter shows the best agreement with the reference solution. The capabilities and potential limitations of this new approach will be further scrutinized in future work.
Acknowledgements
SR has been supported by the Swedish Research Council (VR) under grant number 2020-05044, by the research environment grant “Gravitational Radiation and Electromagnetic Astrophysical Transients” (GREAT) funded by the Swedish Research Council (VR) under Dnr 2016-06012, by the Knut and Alice Wallenberg Foundation under grant Dnr. KAW 2019.0112, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2121 “Quantum Universe” - 390833306 and by the European Research Council (ERC) Advanced Grant INSPIRATION under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101053985).
The simulations for this paper have been performed on the facilities of North-German Supercomputing Alliance (HLRN), and at the SUNRISE HPC facility supported by the Technical Division at the Department of Physics, Stockholm University, and on the HUMMEL2 cluster funded by the Deutsche Forschungsgemeinschaft (498394658). Special thanks go to Mikica Kocic (SU), Thomas Orgis and Hinnerk Stueben (both UHH) for their excellent support.
Some plots have been produced with the software splash (Price, 2007).
Data availability
The data underlying this article will be shared on reasonable request to the corresponding author.